Solve for x: 73 square root of 2x + 5

1
Expert's answer
2021-03-22T09:55:51-0400

73 = "\sqrt{2x + 5}"

"\sqrt{2x + 5}" = 343

("\sqrt{2x + 5}" )2 = 3432

2x + 5 = 117649

2x = 117644

x = 58822

The answer is x = 58822